Chuck, a single taxpayer, earns $85,000 irn taxable income and $10,000 in interest from an investment in City of Heflin bonds. Using the U.S. tax rate schedule, how much federal tax will he owe? What is his average tax rate? What is his effective tax rate? What is his current marginal tax rate? If Chuck earns an additional $40,000 of taxable income, what is his marginal tax rate on this income? If Chuck claims an additional $40,000 of deductions, what is his marginal tax rate on this income??
Answer #1

Hi

Let me know in case any issue and query:

As per 2018 Tax rate schedule: In this case, $10000 interest income is tax exempt as per IRS, therefore, tax will be calculated on $76500 only: Federal tax owed by Chuck: (14089.50) + (85000-82500)*24%- 0 Solution: Answers are highlighted in yellow: a) b) Average tax rate (14689.50)/(85000) c) Effective tax rate (14689.50)/(85000+10000) d) Current Marginal tax rate (the last applicable rate in slab) $ 14,689.50 17.28% 15.46% 24.00%24.00% Answer e) Explanation: Marginal tax rate = Change in Tax / Change in Taxable income Tax on old income (14089.50) + (85000-82500)*24% = Tax on New income (14089.50) + (125000-82500)*24% = Change in Tax Change in Taxable income Marginal Tax Rate 14,689.50 $ 24,289.50 $ 9,600.00 S 40,000.00 24.00% Answer f) Explanation: 22.13% Marginal tax rate = Change in Tax / Change in Taxable income Tax on old income (14089.50) + (85000-82500)*24% Tax on New income (4453.50) + (45000-38700)*22% = Change in Tax Change in Taxable income Marginal Tax Rate $ 14,689.50 $ 5,839.50 S -8,850.00 $-40,000.00 22.13%